Solve: x+2√(x-8)=0
Possible answers are-
a. 4,16
b.-4,2
c.4
d.16

Answer by edjones(8007)

x+2√(x-8)=0
x=-2√(x-8)
x^2=4(x-8)
x^2=4x-32
x^2-4x+30=0
The answer is in complex numbers and none of the 4 choices is correct.